OpenStudy (anonymous):

What is the solution set of this equation? -3|x + 5| + 6 = -15 A. {-12,-2} B. {-12,2} C. {-2,12} D. {2,12}

OpenStudy (anonymous):

B, -12 or 2. -12+5 = |-7| 7x-3 = -21 -21 + 6 = -15 or... 2+5 = 7, 7x-3 = -21 -21+6 = -15